OpenStudy (anonymous):

solve for x log5x+log3=0 please explain

OpenStudy (anonymous):

minus log3 on each side then divide by 5

OpenStudy (anonymous):

multiply e both sides

OpenStudy (angela210793):

log5x*3=log1 15x=1 x=15

OpenStudy (angela210793):

loga+logb=log(a*b)

OpenStudy (anonymous):

\[e ^{\log(5x)+\log(3)}=5x(3) =15x\] \[e^0 = 1\] 15x=1 x = 1/15

OpenStudy (angela210793):

yea sorry X=1/15 :)

OpenStudy (angela210793):

There was no need to use e :)

OpenStudy (anonymous):

true